OpenStudy (anonymous):

the length of pattys garden is 4 meters greater than 3 times the width. The perimeter of her garden is 72 meters. What are the dimensions of pattys garden?

OpenStudy (anonymous):

We start: \[ l=4+3w\\ 2w+2l=72 \]So, we say: \[ 2w+2(4+3w)=72\implies\\ 2w+8+6w=72\implies\\ 8w+8=72\implies\\ 8w=64\implies\\ w=8 \]Since: \[ l=4+3w\implies\\ l=4+3(8)\implies\\ l=28 \]And we are done.

OpenStudy (anonymous):

well that was simple enough when you said it!!! Thanks, i get it!

OpenStudy (anonymous):

Sure thing